# BPE: Behavioral Profiling Ensemble

### v1

@@ 6 Discussion and Future Work
    Construction of Behavioral Profiles: Currently, we construct the behavioral profile through a simple noise injection strategy on the training set. This is a relatively rudimentary implementation, as the primary focus of this work is to highlight the shift in ensemble perspective rather than to pursue the ultimate performance limit. Therefore, developing more sophisticated methods for constructing behavioral profiles will be key to raising the performance ceiling. A distinct advantage of the BPE framework over methods like DES and Stacking is that it does not require a validation set. This characteristic allows for various data augmentation techniques on the training set without the need for ground truth labels, thereby mitigating the limitations imposed by small-sample environments compared to traditional approaches.

### v2

@@ 7.3 Limitations and future works
    Construction of Behavioral Profiles: Currently, we construct the behavioral profile through a simple noise injection strategy on the training set. This serves as a foundational baseline implementation, as the primary focus of this work is to highlight the shift in ensemble perspective rather than to pursue the ultimate performance limit. Therefore, developing more sophisticated methods for constructing behavioral profiles will be key to raising the performance ceiling. A distinct advantage of the BPE framework over methods like DES and Stacking is that it does not require a validation set. This characteristic allows for various data augmentation techniques on the training set without the need for ground truth labels, thereby mitigating the limitations imposed by small-sample environments compared to traditional approaches.